The following transactions of Hector Limited took place during January 2022 in respect of component J used in production:

  • Stock of component J on hand - 150 units @ R24 per unit
  • Issued to production - 110 units
  • Purchased from supplier - 200 units @ R25 per unit
  • Returned to supplier (purchased on 12 January 2022) - 30 units
  • Issued to production - 170 units
  • Purchased from supplier - 140 units @ R26 per unit
  • Issued to production - 130 units


Calculate the value of closing inventory on 31 January 2022 using the:

  1. First-in-first-out (FIFO) method
  2. Weighted Average Cost Method







1
Expert's answer
2022-03-28T15:21:06-0400

FIFO method

Initially there was 150 units @24. 110 units were issued to production so the balance is (150-110)= 40 units. After purchasing 200 units, the new balance is (200+40)= 240 units.30 units were returned to supplier so the new balance is (240-30)=210 units. 170 units were issued to production so the new balance is (210-170)= 40 units. 140 units were then purchased, so the new balance is (140+40)= 180 units. After issuing 130 units to production, the new balance is (180-130)= 50 units

The cost of closing stock is thus (50*26) =1300

So the answer is 1300.


Weighted average method

In this method, we will find the value of the total units purchased and divide it by the total units purchased to find the value of one unit

(150*24)+(200*25)+(140*26)=12240

12240/490=24.98

Since 50 units are remaining their value is (24.98*50)=1249

So the answer is 1249